OpenStudy (anonymous):

Regarding the R formula.. How do you find θ given this: 5sinθ + 2cosθ

ganeshie8 (ganeshie8):

does that equal something ?

OpenStudy (anonymous):

nope well, they asked to find θ for a maximum length

OpenStudy (anonymous):

so basically say length AB AB = 5sinθ + 2cosθ

zepdrix (zepdrix):

R formula? 0_o

zepdrix (zepdrix):

We want to find the angle \(\Large\rm \theta\) which maximizes the length of the curve \(\Large\rm AB\) ?

OpenStudy (anonymous):

AB is a straight line Rformula like asinθ±bcosθ=Rsin(θ∓α)

OpenStudy (anonymous):

Okay maybe the question itself would shed more light  Given that <QOB is a variable angle θ, where 0˚< θ < 90˚, obtain an expression for AB in terms of θ. this i already found, which is AB = 5sinθ + 2cosθ Then express AB in the form of Rsin(θ+α), which is (√25+4)sin(θ+basic angle) =(√29)sin taninverse(2/5) —this is as much as i could simplify??— hence find the value of θ for which AB has a maximum length.